# Hyperbolic recurrent neural network as the first type of non-Euclidean neural quantum state ansatz

### v2

@@ 5.1.1 Experiment details
For the 1D TFIM system, we choose four settings with the number of spins $N$ being 20, 40, 80 and 100. Three different RNN-based real NQS, namely Euclidean RNN, Euclidean GRU and hyperbolic GRU, are used as ansatzes in the VMC experiments for 1D TFIM at each $N$ (see Table 1). The architecture of these ansatzes is the same as that of the real NQS wavefunction described in Section 4.1, which consists of an RNN (either Euclidean or hyperbolic) with the hidden dimension $d_{h}=50$ and a Dense layer with 2 units (representing the 0 or 1 entry of the spin ${\sigma}_{i}$ at site $i$) with the Softmax activation function. The wavefunction is given by Eq.(36) in which the state $|\vec{\sigma}\rangle$ is a vector of length $N$ with each entry assuming a value of either 0 or 1 (spin up or down). The number of spin configuration samples that are generated is kept fixed at 50 for all three ansatzes. This choice is made based on the observation that increasing the number of samples generated does not really improve the results while significantly increasing the training time.
@@
The number of parameters of the Euclidean RNN ansatz is 2752, while the number of parameters of the Euclidean and hyperbolic GRU-based ansatz is 8052. For Euclidean ansatzes, we use Adam optimizer with an exponential decay learning rate while for the hyperbolic ansatz, we use RSGD (Riemannian SGD) with the fixed learning rate of ${\alpha}=10^{-2}$ for hyperbolic parameters and the same Adam with exponentially decaying learning rate for Euclidean parameters. In the hyperbolic GRU ansatz, only the biases ($b_{h},b_{r},b_{z})$ are hyperbolic, while the weight matrices are Euclidean. This configuration is the same as the one used in [30], and is kept unchanged throughout this work.
At each iteration of the training process, a neural network ansatz generates 50 samples (spin configurations) whose mean energy is used to approximate the true ground state energy of the 1D TFIM. Convergence is reached when there is no noticeable change in the mean energy and the energy variance of all the samples is smaller than a certain tolerance threshold. This specific values of the tolerance threshold vary for different VMC experiments, and can range from as low as 0.05 to 1101010In other settings such as 2D TFIM, 1D $J_{1}J_{2}$ models, the variance tolerance can be much higher. For all three types of ansatzes at all $N$, convergence is reached very quickly at around 100 epochs with the total number of training steps set at 120 epochs. Even with this very small number of epochs, for the hyperbolic GRU to complete the training loop, it takes around 1.5 hours for $N=20$ and up to 10 hours for $N=100$ on an average laptop111111a average laptop at the time of writing is one with around 8-16Gb with no GPU. On the other hand, Euclidean RNN and GRU take around 0.1 hours for $N=20$ and around 1.5 hours for $N=100$ on the same hardware. During the training process, only the best models are saved, i.e. the training loop might continue until the specified number of iterations is reached but unless a better result is obtained, the final model is the last saved model, e.g. if the best model is the one at epoch 65, training still carries on until epoch 120 but no new model is saved after epoch 65 unless there is a better model at the $i^{th}$ epoch where $12065<i<$ 120. Unless otherwise stated, these training details (include the type of optimizers, the hardware resource, the number of generated samples used in training, the selected saving of best model) remain the same throughout this work in all VMC experiments.
@@ 5.1.2 Results
In Table 2, we list the training results recorded directly from the training process, while in Table 3, we list the inference results for the VMC experiments for the 1D TFIM at $N=20,40,80,100$. The inference results are obtained by using the trained models to generate new spin configurations to calculate the new NQS wavefunction. The convergence plots for the mean energy and energy variance recorded from the training process are shown in Fig.11 and Fig.12 in Appendix A.1.
@@
Table 2: VMC training results for 1D TFIM Hamiltonian with various system sizes $N=20,40,80,100$ for $J=1$ and $B=1$. These results are recorded directly from the training process and correspond to the parameters of the best saved model. For each ansatz, the mean $E$ is listed first, together with the variance (in brackets). The best training result for the converged mean energy at each $N$ is noted in bold.
Ansatz $N=20$ $N=40$ $N=80$ $N=100$
eRNN-50-s50 -25.0802 -50.5518 -101.3626 -126.8852
(0.074) (0.248) (0.331) (0.141)
eGRU-50-s50 -25.0713 -50.5922 -101.4369 -126.8455
(0.068) (0.220) (0.160) (0.249)
hGRU-50-s50 -25.0725 -50.6399 -101.4554 -126.7369
(0.055) (0.149) (0.481) (0.245)
Exact (DMRG) -25.1078 -50.5694 -101.4974 -126.9619
Table 3: VMC inference results for 1D TFIM Hamiltonian with various system sizes $N=20,40,80,100$ for $J=1$ and $B=1$. For each ansatz, the mean $E$ is listed first, together with the variance (in brackets). These results are obtained from the inference process in which the trained parameters are loaded into new models and these models are used to calculate the mean energy. The best inference result for the converged mean energy at each $N$ is noted in bold.
Figure 4: VMC experiments for 1D TFIM with $N=20,40,80$ and 100 spins (from top to bottom row, left to right): Comparisons of the performances of the variants eRNN-50-s50, eGRU-50-s50, hypGRU-50-s50 (listed in Table 1) - abbreviated as eRNN, eGRU and hGRU on the horizontal axis of each subfigure - during training and inference processes. In each subfigure, the mean energy from training and inference for each NQS ansatz is shown as a dot with error bar (representing the standard error).
The results listed in Table 2 and Table 3 are graphically presented in Fig.4, from which one can see immediately the performances of the three types of NQS ansatzes with respect to one another. We note that it is not out of the ordinary that the performances during training and inference of the same ansatz differ.
  1. 1.
    For $N=20$, the best NQS ansatz is eRNN-50-s50 (Euclidean RNN) from inference and eGRU-50-s50 (Euclidean GRU) from training.
  2. 2.
    For $N=40$, the best NQS ansatz is eRNN-50-s50 (Euclidean RNN) from inference and hGRU-50-s50 (hyperbolic GRU) from training.
  3. 3.
    For $N=80$, the best NQS ansatz is hGRU-50-s50 (hyperbolic GRU) from inference and eGRU-50-s50 (Euclidean GRU) from training.
  4. 4.
    For $N=100$, the best NQS ansatz is eRNN-50-s50 (Euclidean RNN) from both inference and training. For the training process, the three NQS ansatzes have very close performances among one another.
@@ 5.2.1 Experiment details
For the 2D TFIM, we will perform the VMC experiments using three different NQS ansatzes: one-dimensional Euclidean GRU, one-dimensional hyperbolic GRU and two-dimensional Euclidean RNN, the details (including the names and number of parameters) of which are included in Table 4. The size $(N_{x},N_{y})$ of the square lattice in the 2D TFIM is chosen to be (5,5), (7,7), (8,8) and (9,9) [^12] Unfortunately, we do not have enough computing resources to carry out the VMC experiments for the case of $(N_{x},N_{y})\geq(10,10)$ corresponding to systems with 25, 49, 64 and 81 spins. All neural network ansatzes are real NQS wavefunction (described in Section 4.1) whose architecture comprises an RNN unit (either Euclidean GRU, hyperbolic GRU or 2D Euclidean RNN) with the hidden dimension $d_{h}=50$, and a Dense layer with the Softmax activation function.
@@
where $f$ is a nonlinear activation function, the input $\vec{x}_{t}=(\vec{x}^{0}_{t},\vec{x}^{1}_{t})$ and hidden state $\vec{h}_{t}=(\vec{h}^{0}_{t},\vec{h}^{1}_{t})$ at timestep $t$ are now a 2D vectors comprising the horizontal and vertical components. Correspondingly, there is a doubling of the ususal weight matrices $U$ and $W$ to $U_{h},U_{v}$, $W_{h},W_{v}$. For a 2D spin configuration $\mathbf{{\sigma}}$ of dimension $(N_{x},N_{y}$), an entry $s_{i,j}$ is autoregressively generated based on the two spins and two hidden states from its previously generated horizontal and vertical neighbors. For the 1D Euclidean/hyperbolic GRU, during the sampling process, the site ${\sigma}_{i,j}$ is autoregressively generated based on only the previously generated site in a raster manner shown in Fig.5. At each time step $t$ during the training phase, the 2D Euclidean RNN NQS ansatz generates 50 two-dimensional samples, each of which is a (2D) spin configuration (samples) of size $(N_{x},N_{y}$), while the 1D Euclidean/hyperbolic GRU ansatz generates 50 one-dimensional samples, each of which is a (1D) spin configuration of length $N_{x}\times N_{y}$. The computation of the energy for the 1D ansatz requires a reshaping of the 1D samples of length $N_{x}\times N_{y}$ into 2D samples of size $(N_{x},N_{y})$.
 11\,1\,1 22\,2\,2 33\,3\,3 44\,4\,4 55\,5\,5 66\,6\,6 77\,7\,7 88\,8\,8 99\,9\,910101010111111111212121213131313141414141515151516161616
 11\,1\,1 22\,2\,2 33\,3\,3 44\,4\,4 88\,8\,8 77\,7\,7 66\,6\,6 55\,5\,5 99\,9\,910101010111111111212121213131313141414141515151516161616
Other training details are largely the same as described in the 1D TFIM, including the hardware configuration, the fixed (50) number of samples generated during training, the optimizer types (Adam for Euclidean NQS, with the learning rate of $10^{-2}$ with exponential decay every 100 epochs, and a combination of Adam+RSGD for hyperbolic NQS where the learning rate setting is the same for Adam in the Euclidean case while for RSGD, it is a fixed learning rate of $10^{-2}$ throughout) and the setup in which only the best model is saved during the training phase. For $(N_{x},N_{y})=(5,5),(7,7)$, with the number of training epochs being 450, 1D Euclidean GRU takes less than 2 hours, while 1D hyperbolic GRU takes up to 14 hours to finish the training. For $(N_{x},N_{y})=(8,8),(9,9)$, with the number of training epochs reduced 350 due to the significantly longer running time required, 1D Euclidean GRU takes around 4 hours, while 1D hyperbolic GRU can take up to around 30 hours to finish the training. For the lighter 2D Euclidean RNN that reaches convergence faster than the 1D ansatzes, the number of epochs varies and is 200 for $(N_{x},N_{y})=(5,5)$, 300 for $(N_{x},N_{y})=(7,7)$ and 350 for $N(_{x},N_{y})=(8,8),(9,9)$. The training time of 2D Euclidean RNN ranges from 0.2 hour for $(N_{x},N_{y})=(5,5)$ to 3.5 hours for $(N_{x},N_{y})=(9,9)$.
@@ 5.2.2 Results
The training and inference results for the VMC experiments involving the 2D TFIM setting are listed in Table 5 and Table 6. The convergence curve plots for the mean energy and energy variance are shown in Fig.13 and Fig.14 in Appendix A.2.
@@
Table 5: VMC training results for 2D TFIM Hamiltonian with various system sizes $(N_{x},N_{y})$ for $J=1$ and $B=3.0$. These results are recorded directly from the training process and correspond to the parameters of the best saved model. For each ansatz, the mean $E$ is listed first, together with the variance (in brackets). The best and second best results for the converged mean energy for each $(N_{x},N_{y})$ case are noted in bold and typewriter font, respectively. For this 2D TFIM setting, we care more about the second best than the best result, since the best result in each case is expected to be 2D Euclidean RNN.
1d-eGRU-50-s50 -78.1900 -153.3293 -200.0871 -255.4191
(5.300) (14.429) (26.120) (28.418)
1d-hGRU-50-s50 -78.3403 -154.4542 -200.5193 -255.6020
(4.480) (7.054) (28.783) (19.538)
2d-eRNN-50-s50 -78.6763 -154.6495 -202.3759 -256.7026
(0.390) (0.448 ) (1.083) (0.881)
DMRG -78.6857 -154.8463 -202.5077 -256.5535
@@
Table 6: VMC inference results for 2D TFIM Hamiltonian with various system sizes $(N_{x},N_{y})$ for $J=1$ and $B=3.0$. These results were obtained from the inference process in which the trained model is used to generate new spin configurations whose mean energies are then calculated. For each ansatz, the mean $E$ is listed first, together with the variance (in brackets). The best and second best results for the converged mean energy for each $(N_{x},N_{y})$ case are noted in bold and typewriter font, respectively. For this 2D TFIM setting, we care more about the second best than the best result, since the best result in each case is expected to be 2D Euclidean RNN.
From Table 5, Table 6 and Fig.6, it is clear that 2D Euclidean RNN is the best performing ansatz in all four cases for both training and inference, since the architecture of the RNN part of this NQS was specially created for the 2D setting of TFIM. It is thus within our expectation that 2D RNN will outperform both 1D Euclidean and hyperbolic GRU. What we are really interested in, and would like to find out, however, is whether 1D hyperbolic GRU will outperform 1D Euclidean GRU. Regarding this aspect, we observe that in the training process, hyperbolic GRU outperform Euclidean GRU 3 out of 4 times when $(N_{x},N_{y})=(5,5),(7,7),(9,9)$, while in inference process, hyperbolic GRU outperform Euclidean GRU 4 out of 4 times (see Fig.6). While for $(N_{x},N_{y})=(5,5),(7,7)$, it is clear that hyperbolic GRU definitively outperform Euclidean GRU for in both training and inference, the situation is not clear-cut for $(N_{x},N_{y})=(8,8),(9,9)$. When $(N_{x},N_{y})=(9,9)$ Euclidean and hyperbolic GRU are almost comparable in performance during training, while for $N_{x},N_{y})=(8,8)$, their inference results are almost comparable, although for both cases, hyperbolic GRU still emerged as the slightly better variant. Also, from the converge curve for the mean energy, Fig.13, we note that the cyan curve corresponding to 1D hyperbolic GRU is always closer, with narrower range of fluctuations, to the DMRG line than the blue curve corresponding to 1D Euclidean GRU, even for the case of $(N_{x},N_{y})=(8,8)$, when the final saved 1D hyperbolic GRU model does not yield the result as good as that of 1D Euclidean GRU (see the first subfigure on the second row of Fig.13). While more experiments are needed, especially for larger $(N_{x},N_{y})$ sizes, from the results of these experiments, it can be seen that 1D hyperbolic GRU is likely to outperform 1D Euclidean GRU.
@@
Figure 6: Comparisons of the performances (during training and inference) of 2D Euclidean RNN and 1D Euclidean/hyperbolic GRU ansatzes listed in Table 4 for the 2D TFIM with different ($N_{x},N_{y}$) lattices - clockwise from top: $(N_{x},N_{y})=(5,5)$, (7,7), (9,9), (8,8). Each dot with error bar represents the mean $E$ value of an NQS ansatz. For all cases, 2D Euclidean RNN is the best ansatz. The second best ansatz is 1D hyperbolic GRU for all cases, except for the training result of $(N_{x},N_{y})=(8,8)$.
[…]
@@
This result is very interesting since it might be indicative of the role played by the hyperbolic geometry of the NQS ansatz in Hamiltonian systems where there exist some sort of hierarchical structure. In the context of natural language processing (NLP) when the hyperbolic RNN/GRU was introduced in the work of [30], the authors noted a definitive and clear outperformance of hyperbolic RNN/GRU over their Euclidean version when the underlying data exhibit a hierarchical, tree-like structure. In the present 2D TFIM setting, there exists a hierarchy in the interaction structure of 2D Hamiltonian when 1D NQS is used. This is due to the artificial rearrangement (or folding) of the 1D samples generated by the 1D NQS into 2D samples during the calculations of their local energies (see Fig.7). This process of rearrangement creates additional interactions between neighbors at site $i$ and site $i+N$, which were originally faraway in the 1D horizontal chain but are brought close together to form the vertical dimension in the 2D square lattice. Concretely, let us consider a chain of length $N^{2}=(N-1)N+N$ spins (see Subfigure (a) of Fig.7) folded into a square lattice of size ($N,N$) in Subfigure (b) of Fig.7. In 1D when only the next nearest neighbor interactions are counted, site $i$ and site $i+N$ have no interaction, but when arranged in a 2D square lattice, these sites are each other next nearest neighbor vertically, which means their interactions are counted in the local energy calculations. Together with the original 1D next nearest neighbor interactions $\langle i,i+1\rangle$, these additional interactions $\langle i,i+N\rangle$ actually form a hierarchy of interactions [^13] Note that alongside the gained interaction $\langle i,i+N\rangle$, we also lose some next neareast neighbor interactions $\langle i,i+1\rangle$ when $i=kN$, with $k\in\mathbb{N}$, i.e. when $i$ is multiple of $N$, and it is this hierarchy that might have played a part in the observation of 1D hyperbolic GRU outperforming 1D Euclidean GRU in 2D TFIM setting.

### v3

@@ 5.1.1 Experiment details
For the 1D TFIM system, we choose four settings with the number of spins $N$ being 20, 40, 80 and 100. Three different RNN-based real NQS, namely Euclidean RNN, Euclidean GRU and hyperbolic GRU, are used as ansatzes in the VMC experiments for 1D TFIM at each $N$ (see Table 1). The architecture of these ansatzes is the same as that of the real NQS wavefunction described in Section 4.1, which consists of an RNN (either Euclidean or hyperbolic) with the hidden dimension $d_{h}=50$ and a Dense layer with 2 units (representing the 0 or 1 entry of the spin ${\sigma}_{i}$ at site $i$) with the Softmax activation function. The wavefunction is given by Eq.(36) in which the state $|\vec{\sigma}\rangle$ is a vector of length $N$ with each entry assuming a value of either 0 or 1 (spin up or down).
@@
The number of parameters of the Euclidean RNN ansatz is 2752, while the number of parameters of the Euclidean and hyperbolic GRU-based ansatz is 8052. For all three types of ansatzes at all $N$, convergence is reached very quickly at around 100 epochs with the total number of training steps set at 120 epochs. Even with this very small number of epochs, for the hyperbolic GRU to complete the training loop, it takes more than 1.5 hours for $N=20$ and up to 20 hours for $N=100$ on an average laptop141414a average laptop at the time of writing is one with around 8-16Gb with no GPU. On the other hand, Euclidean RNN and GRU take around 0.1 hours for $N=20$ and around 1.5 hours for $N=100$ on the same hardware.
@@ 5.1.2 Results
In Table 2, we list the inference results, obtained by using the trained models to generate new spin configurations to calculate the new NQS wavefunction, for the VMC experiments for the 1D TFIM at $N=20,40,80,100$. The number of samples used for inference is $10^{4}$. The convergence plots for the mean energy recorded from the training process are shown in Fig.11 in Appendix B.1.
@@
Table 2: VMC inference results (using 10000 samples) for 1D TFIM Hamiltonian with various system sizes $N=20,40,80,100$ for $J=1$ and $B=1$. For each ansatz, the mean $E$ is listed first, together with the corresponding standard error directly below (in brackets). These results are obtained from the inference process in which the trained parameters are loaded into new models and these models are used to calculate the mean energy. The best inference result for the converged mean energy at each NN is noted in bold.
Figure 4: VMC experiments for 1D TFIM with $N=20,40,80$ and 100 spins (from top to bottom row, left to right): Comparisons of the performances of the variants eRNN-50-s50, eGRU-50-s50, hypGRU-50-s50 (listed in Table 1) - abbreviated as eRNN, eGRU and hGRU on the horizontal axis of each subfigure. In each subfigure, the mean energy for each NQS ansatz is shown as a dot with an error bar (representing the standard error).
The results listed in Table 2 are graphically presented in Fig.4, from which one can see immediately the performances of the three types of NQS ansatzes with respect to one another. In all cases, Euclidean GRU and hyperbolic GRU NQS outperformed Euclidean RNN NQS. However, when it comes to the comparing the performances of Euclidean and hyperbolic GRUs, the situation is not as clear-cut.
  • For $N=20,100$, the best NQS ansatz is hGRU-50-s50, with the second best one being eGRU-50-s50.
  • For $N=40,80$, the best NQS ansatz is eGRU-50-s50, with the second best one being hGRU-50-s50.
@@ 5.2.1 Experiment details
For the 2D TFIM, we will perform the VMC experiments using three different NQS ansatzes: one-dimensional Euclidean GRU, one-dimensional hyperbolic GRU and two-dimensional Euclidean RNN (see Appendix A for the detailed description of the 2D RNN neural network and the sample generation process involving both 1D and 2D NQS), the details (including the names and number of parameters) of which are included in Table 3. Since the 2D Euclidean RNN NQS was specially designed by taking into account the 2D geometry of the TFIM physical system, it is our expectation that 2D RNN will outperform both 1D Euclidean and hyperbolic GRU. The goal of this experiment, therefore, is not to determine the best, but rather the second-best performing NQS. In other words, what we are really interested in, and would like to find out is the performance of 1D hyperbolic GRU with respect to 1D Euclidean GRU. The inclusion of the 2D RNN NQS was made to provide an overall benchmark in the 2D TFIM case.
The size $(N_{x},N_{y})$ of the square lattice in the 2D TFIM is chosen to be (5,5), (7,7), (8,8) and (9,9) [^15] Unfortunately, we do not have enough computing resources to carry out the VMC experiments for the case of $(N_{x},N_{y})\geq(10,10).$ corresponding to systems with 25, 49, 64 and 81 spins. All neural network ansatzes are real NQS wavefunction (described in Section 4.1) whose architecture comprises an RNN unit (either Euclidean GRU, hyperbolic GRU or 2D Euclidean RNN) with the hidden dimension $d_{h}=50$, and a Dense layer with the Softmax activation function.
@@
Other training details are largely the same as described in the 1D TFIM, and will not be repeated here. For $(N_{x},N_{y})=(5,5),(7,7)$, with the number of training epochs being 450, 1D Euclidean GRU takes less than 2 hours, while 1D hyperbolic GRU takes up to 14 hours to finish the training. For $(N_{x},N_{y})=(8,8),(9,9)$, with the number of training epochs reduced to 350 due to the significantly longer running time required, 1D Euclidean GRU took around 4 hours, while 1D hyperbolic GRU could take up to around 50 hours to finish the training. For the lighter 2D Euclidean RNN that reached convergence faster than the 1D ansatzes, the numbers of epochs varied and were 200 for $(N_{x},N_{y})=(5,5)$, 300 for $(N_{x},N_{y})=(7,7)$ and 350 for $(N_{x},N_{y})=(8,8),(9,9)$. The training times of 2D Euclidean RNN ranged from 0.2 hour for $(N_{x},N_{y})=(5,5)$ to 3.5 hours for $(N_{x},N_{y})=(9,9)$.
@@ 5.2.2 Results
The inference results (using 10000 samples) for the VMC experiments involving the 2D TFIM setting are listed in Table 4. The convergence curve plots for the mean energy are shown in Fig.12 in Appendix B.2.
@@
Table 4: VMC inference results for 2D TFIM Hamiltonian with various system sizes $(N_{x},N_{y})$ for $J=1$ and $B=3.0$. These results were obtained from the inference process in which the trained model is used to generate new spin configurations whose mean energies are then calculated. For each ansatz, the mean $E$ is listed first, together with the corresponding standard error directly below (in brackets). The best results for each $(N_{x},N_{y})$ case are noted in typewriter font while the second best results are noted in bold. We emphasize that for this experiment, it is the second best results that we are interested in (since the best result in each case is expected to be 2D Euclidean RNN)
From Table 4, as expected, 2D Euclidean RNN is the best performing ansatz in all four cases, since the architecture of the RNN part of this NQS was specially created for the 2D setting of TFIM. Again, we stress that the goal of this experiment is not to determine the best, but rather, the second-best performing NQS. We are solely interested in determining the relative performance of 1D hyperbolic GRU with respect to 1D Euclidean GRU in the setting of 2D TFIM, with the inclusion of the 2D RNN NQS ansatz being made only to provide an additional benchmark. Regarding this aspect, we observe that hyperbolic GRU outperformed Euclidean GRU 4 out of 4 times (see Fig.5). While for $(N_{x},N_{y})=(7,7),(8,8),(9,9)$, it is clear that hyperbolic GRU definitively outperformed Euclidean GRU, the situation is not clear-cut for $(N_{x},N_{y})=(5,5)$ when the performances are almost comparable, with hyperbolic GRU emerging as the slightly better ansatz. While more experiments involving larger $(N_{x},N_{y})$ sizes are needed, from the results of these experiments, it can be seen that 1D hyperbolic GRU would likely outperform 1D Euclidean GRU in 2D TFIM experiments.
@@
Figure 5: Comparisons of the performances of 1D Euclidean GRU and 1D hyperbolic GRU ansatzes listed in Table 3 for the 2D TFIM with different ($N_{x},N_{y}$) lattices - clockwise from top: $(N_{x},N_{y})=(5,5),(7,7),(9,9),(8,8)$. Each dot with error bar represents the mean $E$ value of an NQS ansatz. For all cases, 2D Euclidean RNN is the best ansatz and is not included in this comparison, which strictly measures the performances of the 1D Euclidean versus 1D hyperbolic GRU. The second best ansatz is always 1D hyperbolic GRU for all $(N_{x},N_{y})$.
@@
This result is very interesting since it might be indicative of the role played by the hyperbolic geometry of the NQS ansatz in Hamiltonian systems where there exist some sort of hierarchical structure. In the context of natural language processing (NLP) when the hyperbolic RNN/GRU was introduced in the work of [30], the authors noted a definitive and clear outperformance of hyperbolic RNN/GRU over their Euclidean version when the underlying data exhibit a hierarchical, tree-like structure. In the present 2D TFIM setting, there exists a hierarchy in the interaction structure of 2D Hamiltonian when 1D NQS is used.
This is due to the artificial rearrangement (or folding) of the 1D spin chain samples generated by the 1D NQS to model the 2D data during the calculations of their local energies (see Fig.10 in Appendix A). This process of rearrangement takes into account interactions between neighbors at site $i$ and site $i+N$, which were originally faraway in the 1D horizontal chain but are brought close together to form the vertical dimension in the 2D square lattice. Concretely, let us consider a chain of length $N^{2}=(N-1)N+N$ spins (see Subfigure (a) of Fig.10) folded into a square lattice of size ($N,N$) in Subfigure (b) of Fig.10. In 1D when only the next nearest neighbor interactions are counted, site $i$ and site $i+N$ have no interaction, but when arranged in a 2D square lattice, these sites are each other next nearest neighbor vertically, which means their interactions are counted in the local energy calculations. Together with the original 1D next nearest neighbor interactions $\langle i,i+1\rangle$, these additional interactions $\langle i,i+N\rangle$ actually form a hierarchy of interactions [^16] Note that in the folding process of the 1D spin chain to model the 2D data, alongside the ‘gained’ (in the sense that these are included in the Hamiltonian calculations due to the 2D geometry) interaction $\langle i,i+N\rangle$, we also ‘lose’ some next neareast neighbor interactions $\langle i,i+1\rangle$ when $i=kN$, with $k\in\mathbb{N}$, i.e. when $i$ is multiple of $N$., and it is this hierarchy that might have played a part in the observation of 1D hyperbolic GRU outperforming 1D Euclidean GRU in 2D TFIM setting.